In Islamabad city the mean number of power outages is 4 per month. Find the
probability that in a given month,
a. There are exactly 3 outages
b. There are more than 2 outages.
1
Expert's answer
2020-06-16T18:18:00-0400

The number of power outages follows a Poisson Distribution.

a) P(x=3)=(λxe)/x!=(43e-4)/3!=0.195

b)P(x>2)=1-P(x=0)-P(x=1)-P(x=2)=1-(40e-4)/0!-(41e-4)/1!-(42e-4)/2!=1-0.0183-0.07326-0.1465=0.76194